### v1

# QUEST-A: Untrained Filtering with Trained Focusing led to Enhanced Quantum Architectures

Abstract
Quantum architecture search (QAS) represents a fundamental challenge in quantum machine learning, with SOTA methods primarily categorized into training-free and gradient-guided approaches. However, treating QAS solely as either a discrete pruning process or a continuous optimization problem fails to balance accuracy and efficiency. This work decomposes QAS into two alternately solved subproblems: optimal circuit structure retrieval and parameter optimization. Based on this insight, we propose Quantum Untrained-Explored Synergistic Trained Architecture (QUEST-A), which implements rapid architecture pruning through inherent circuit properties and develops focused optimization with parameter reuse strategies. QUEST-A unifies discrete structure search and continuous parameter optimization within an evolutionary framework that integrates rapid pruning and fine-grained optimization. Experiments demonstrate QUEST-A’s superiority over existing methods: enhancing model expressivity in signal representation, maintaining high performance across varying complexities in image classification, and achieving order-of-magnitude precision improvements in variational quantum eigensolver tasks. These results validate QUEST-A’s effectiveness and provide transferable methodologies for QAS.
keywords:
Quantum machine learning , Quantum architecture search , Parameter reuse , Architecture pruning
journal: Knowledge-Based Systems
[…]
[table]capposition=top
[…]
1 Introduction
Quantum Machine Learning (QML) aims to reinterpret and enhance classical machine learning within the framework of quantum computing Biamonte et al. (2017). QML not only inherits the wide applicability of traditional machine learning, encompassing diverse tasks from classification Huang et al. (2021); Wang et al. (2022); Anagolum et al. (2024) and regression to generative modeling Niu et al. (2022); Zoufal et al. (2021) and reinforcement learning Huang et al. (2022), but also incorporates the unique advantages of quantum computing. In the era of Noisy Intermediate-Scale Quantum (NISQ) devices Preskill (2018), QML has garnered significant attention for its potential to solve complex real-world problems.
Currently, Variational Quantum Circuits (VQCs) represent the most widely adopted method in QML Gong et al. (2024). VQCs optimize model performance through iterative optimization of quantum circuit parameters until the quantum circuit can learn and represent the input problem with high accuracy Thanasilp et al. (2023). From this perspective, QML can be abstracted as a architecture optimization problem guided by a loss function and subject to multiple quantum constraints. This perspective, though radical, bridges the gap between classical machine learning and its quantum counterpart. While VQCs provide a powerful framework for quantum machine learning, designing efficient quantum circuits faces significant challenges as problem scales increase. Traditional manual design methods, although highly task-specific, often struggle to handle complex, large-scale problems effectively. Conversely, the adoption of generalized template-based quantum circuit designs, while simplifying the design process, frequently leads to severe performance bottlenecks, particularly when confronted with issues such as barren plateaus McClean et al. (2018).
To address these challenges, Quantum architecture search (QAS) has emerged as a promising solution Wang et al. (2022); Du et al. (2022); Zhang et al. (2022); Wu et al. (2023); He et al. (2024b). QAS ingeniously combines classical machine learning optimization strategies with fundamental quantum physics principles, aiming to automatically discover optimal quantum circuit structures that satisfy learning task requirements under multiple constraints.
The core challenge of QAS lies in searching for optimal quantum circuit architectures that maintain expressive power while ensuring computational feasibility Lu et al. (2023); Ge et al. (2024). Currently, most QAS methods draw inspiration from the Neural Architecture Search (NAS) paradigm and can be broadly categorized into heuristic, differential-based, and sampling-based approaches based on their optimization strategies. Heuristic algorithms Wang et al. (2022); Williams and Gray (1998); Lamata et al. (2018); Yu et al. (2024) address the circuit architecture optimization problem through quadratic modeling and search the solution space according to predefined rules. These methods offer strong interpretability, providing clear insights into the optimization process. However, they suffer from limited scalability and can only probabilistically obtain global optima, often converging to local optima in large-scale problems. As an extension of variational quantum algorithms, differential-based QAS Zhang et al. (2022); Wu et al. (2023); Zhang et al. (2021); Du et al. (2022); Ma et al. (2024) constructs end-to-end differentiable workflows, utilizing gradient-guided searches with predefined supercircuits or modules. This approach effectively leverages gradient information; nevertheless, its search efficiency is constrained by preset structures and optimization techniques, resulting in computationally intensive training processes. Training-free QAS Anagolum et al. (2024); He et al. (2024b), an emerging efficient search mode derived from classical training-free architecture search methods Zhou et al. (2022), employs intrinsic characteristics of quantum circuits as performance proxy indicators to rapidly screen candidate circuit sets. The strength of this approach lies in its search efficiency, but rapid, undirected nature of the search also risks overlooking potentially superior architectural configurations.
Figure 1: Comparison between gradient-based and training-free QAS methods. Output single-qubit states are used to visualize the circuit search process. (a) Continuous gradient-based approach gradually converges to target circuit through gradient guidance. (b) Discrete training-free approach narrows search space through screening indicators to locate target circuit vicinity.
Considering the respective strengths and limitations of the aforementioned QAS methodologies, this research is motivated by the imperative to address the trade-off between expressibility and trainability in QAS. To this end, we propose the QUEST-A, which aims to surmount the inherent limitations of singular paradigms by synergistically combining the advantages of rapid untrained exploration and fine-grained trained optimization.
Specifically, QUEST-A adopts a multi-level search strategy, combining training-free rapid screening with variational quantum circuit fine-tuning optimization. It employs reupload circuits as a template for flexible architecture exploration. During the rapid screening phase, QUEST-A evaluates quantum circuits through complexity and expressivity indicators, which is coupled with subsequent gradient-based optimization to ensure search quality. The multi-level strategy naturally forms an iterative evolutionary process, where weight-sharing strategies accumulate performance advantages across generations, unifying discrete structure search and continuous parameter optimization. The contributions of our work are as follows:
  1. 1.
    Proposed the QUEST-A framework, integrating untrained rapid filtering with trained focusing strategies, overcoming the limitations of existing methods that overly emphasize either efficiency or accuracy, and achieving an optimized balance between search process efficiency and precision.
  2. 2.
    Viewed the natural evolution of the search process as an evolutionary sequence of candidate circuit ensembles, establishing rapid indexing and parameter transfer mechanisms. This approach achieved efficient transmission of advantageous parameters.
  3. 3.
    In numerical simulations and NISQ environment experiments across multiple types of tasks, QUEST-A demonstrated performance superior to state-of-the-art methods. Specifically, it achieved up to $81.8\%$ performance improvement in representation tasks, maintained over $75\%$ high accuracy in $10$-class image classification tasks, and enhanced precision by up to $11$ orders of magnitude in VQE tasks.
2 Related Works
2.1 Quantum circuits for quantum machine learning
VQCs have emerged as the dominant methodology in QML. VQCs comprise modular circuits, encompassing data encoding, trainable variational gates, and measurement operations, thereby forming a quantum-classical hybrid computational process as show in Fig. 2.
Figure 2: Quantum-classical hybrid machine learning framework.
This process initiates with the encoding of classical data into quantum states, followed by the manipulation of these states through variational and entangling gates, and culminates in the conversion of results back to classical data via measurement. The training process of VQCs involves parameter optimization, typically employing gradient descent methods. Gradient computation in quantum circuits necessitates multiple circuit executions, with computational costs scaling linearly with the number of parameters. This characteristic, combined with the noise sensitivity of NISQ devicesPreskill (2018), underscores the significant impact of circuit structure on QML performance.
The influence of quantum circuit structure on learning efficacy is primarily manifested in three aspects: expressive power, training efficiency, and hardware compatibility. The intrinsic expressibility of a circuit is determined by the type, quantity, and arrangement of gates, independent of the parameter optimization process, and represents an inherent property of quantum circuits Sim et al. (2019). Training efficiency reflects the trade-off between expressibility and the number of trainable parameters -- complex circuits do not necessarily enhance expressibility but often increase learning costs. In the NISQ era, circuit design can also consider hardware connectivity constraints and noise characteristics.
2.2 Quantum Architecture Search
QAS is an optimization method aimed at finding optimal quantum circuit structures and parameters within a large-scale search space. The complexity of QAS can be quantified by analyzing its search space: as described in Fig. 2, consider a quantum circuit model comprising $n$ qubits and m layers, where each layer can select from a set of quantum gates $G$ with cardinality $k=|G|$. Under this model, the search space size reaches $O(k^{n×m})$ Ge et al. (2024). This exponential growth indicates that the computational complexity of QAS increases significantly as the number of qubits and circuit depth increase. Furthermore, QAS involves both discrete structure search and continuous parameter optimization Zhang et al. (2022), further augmenting the problem’s complexity.
Although the precise computational complexity classification of QAS remains an open question, there are substantial grounds to conjecture that its complexity is at least comparable to NP-hard problems Wu et al. (2023); Díez-Valle et al. (2021). The QAS problem can be formally described as follows: Let the candidate quantum gates be represented by matrix $M\in G^{n\times m}$, define $U_{ij}=\sigma(M_{ij})$, where $\sigma$ maps the quantum gate on the $i$-th qubit and $j$-th layer to a $2^{n}\times 2^{n}$ unitary matrix. The unitary matrix corresponding to the entire quantum circuit can be expressed as follow
$U=\prod_{j=1}^{m}\prod_{i=1}^{n}U_{ij}=\prod_{j=1}^{m}\prod_{i=1}^{n}\sigma(M_{ij}),$ (1)
Based on this, the optimization objective of QAS can be formalized as $\min_{M\in G^{n\times m}}L(U,D),$ where $L$ denotes the loss function, and $D$ represents the dataset or target unitary matrix for the specific problem. The core of this optimization problem lies in identifying the quantum circuit structure M and its parameters that minimize the corresponding loss function for the specific problem within the $O(k^{n×m})$-scale search space. It is noteworthy that $G^{n\times m}$ represents the set of possible quantum gate combinations, the selection of which directly impacts the size and complexity of the search space.
Meticulously designed quantum circuit templates can reduce the value of $k$, thereby diminishing the scale of the search space and making $O(k^{n×m})$ space more tractable. Consequently, the selection of high-quality quantum circuit templates as starting points for the search process becomes paramount.
2.3 Ansatz for QAS
In quantum computing, an ansatz refers to a parameterized quantum circuit structure. In QAS, ansatzes manifest as quantum circuit templates, defining the fundamental structure and characteristics of the search space.
The cascading CNOT circuit is currently the most commonly used ansatz McClean et al. (2018). By constructing complex entangled states, it establishes global correlations between qubits and serves as a prevalent template due to its universality. However, this structure assigns equal weights to all interactions, potentially obscuring feature importance Wang et al. (2022), and the complex entanglement patterns increase the difficulty of parameter optimization. The advancement of explainability of quantum computing Canatar et al. (2022); Cerezo et al. (2021); McClean et al. (2018) has given rise to various novel ansatzes. Hardware-efficient ansatzes Anagolum et al. (2024) consider the specific structure and gate sets of quantum hardware, while tensor network-inspired ansatzes Lazzarin et al. (2022) leverage theoretical insights from classical tensor networks.
Figure 3: Hybrid machine learning framework based on re-uploading quantum circuits. Data embedding becomes a repeatable and optimizable part of the variational circuit.
As an emerging quantum circuit ansatz, re-uploading circuits Pérez-Salinas et al. (2020) alternate data encoding and parameter variation processes, enabling multiple utilizations of input data and thus enhancing the circuit’s expressivity. This design establishes a complex, data-oriented non-unitary process that can approximate any continuous function Zhao et al. (2024), while providing increased flexibility and controllability in the search space. More importantly, by incorporating data encoding into the optimization scope, re-uploading circuits serve as an ideal template for QAS to potentially uncover superior problem representations Ono et al. (2023).
3 Methodology
3.1 Overview of QUEST-A
Viewing QAS as a holistic optimization problem is challenging due to its large search space, suggesting the need for a decomposed optimization approach. A key insight of this study is that decomposing the QAS problem into two subproblems -- optimal circuit structure retrieval and optimal parameter optimization -- and solving these alternately is considerably simpler than addressing the overall optimization problem directly.
Based on this approach, QUEST-A optimizes quantum circuit structures through the synergistic effect of unsupervised exploration and supervised training. As shown in Fig. 4, the QUEST-A framework consists of three core components corresponding to untrained exploration, supervised training, and synergistic search strategies:
  1. 1.
[…]
  2. 2.
    Focused Architecture Optimization: QUEST-A applies variational quantum algorithms for fine-grained parameter optimization on pruned candidate structures. This process combines stochastic gradient descent for efficient parameter tuning. The introduced parameter reuse strategy transfers and utilizes learned parameter information across different structures through weight sharing.
  3. 3.
    Evolutionary Synthesis of Search Strategies: QUEST-A employs evolutionary algorithms as a meta-optimization framework, organically integrating rapid pruning and fine-tuned optimization processes. Population initialization corresponds to candidate generation in rapid architecture pruning, fitness evaluation combines untrained metrics and supervised training performance, selection operations are based on multi-objective optimization criteria, while crossover and mutation operations generate new candidate structures by manipulating quantum gate sequences and connection topologies.
These three coupled components form a closed-loop iterative optimization system. By alternating coarse-grained rapid pruning and fine-grained precise optimization within the evolutionary framework, QUEST-A achieves efficient, adaptive exploration of the quantum architecture space.
Figure 4: QUEST-A framework.
3.2 Rapid Architecture Pruning
QAS can be considered an optimization problem, where appropriate reduction of the optimization space facilitates the optimization process. Rapid architecture pruning primarily focuses on discrete quantum circuit structures, thus further simplifying the problem into a discrete database retrieval task. Consequently, establishing an appropriate indexing mechanism is crucial to expedite the search process. Traditional supercircuit methods Wang et al. (2022); Du et al. (2022), typically construct and store the entire supercircuit at once as indexing , resulting in substantial memory overhead and training costs.
In contrast, QUEST-A adopts a dynamic indexing strategy, effectively overcoming these limitations. Formally, given the number of qubits $N$, maximum circuit depth $L$, and a set of available quantum gates $G(|G|=Q)$, the indexer $A$ for the search space $S$ can be defined as $A:{0,1,...,Q-1}^{N×L}\textrightarrow S$, where ${0,1,...,Q-1}^{N×L}$ represents the $N×L$-dimensional discrete index space. In the $t$-th iteration, the sampled specific quantum circuit structure $a^{t}$ can be expressed as $a^{t}=A(i^{t}),i^{t}\in{0,1,...,Q-1}^{N×L}$. This dynamic indexing method enhances memory efficiency by avoiding the need to store the entire supercircuit structure, while also improving computational efficiency by generating and evaluating specific structures only when necessary.
Furthermore, through the integration of re-uploading circuits in Section  2.3, QUEST-A enables the search for optimal data embedding methods Pérez-Salinas et al. (2020). Specifically, this design partitions the overall encoding task into individual circuit modules, optimizing the embedding process for each module through QAS.
Building upon the QUEST-A framework, the intrinsic properties of quantum circuits are further utilized to optimize the search process. These properties, directly related to the structural complexity and potential expressivity of the circuits, enable rapid identification and exclusion of candidate circuits with evident deficiencies in circuit scale and expressivity, thereby narrowing the search space He et al. (2024b). QUEST-A employs two complementary evaluation methods: topological analysis based on Directed Acyclic Graphs (DAGs) and expressivity assessment using Kullback-Leibler (KL) divergence.
Quantum circuits can be represented as DAGs Holmes et al. (2022), where nodes represent quantum gate operations and edges represent qubit evolution paths. Based on this representation, the number of paths from input to output in the DAG is used as a proxy indicator for circuit complexity. For a given quantum circuit c, its path count is defined as $P(c)$. As demonstrated by the empirical study in Fig. 5, $P(c)$ generally exhibits a positive correlation with the number of CNOT gates and circuit depth. This correlation highlights the effectiveness of $P(c)$ as an indicator of circuit complexity.
Figure 5: Relationship between DAG-path count and circuit complexity for $500$ random quantum circuits. As the DAG-path count increases, both circuit depth and number of CNOT show logarithmic growth, indicating a strong correlation between DAG-path count and circuit scale.
The use of $P(c)$ aims to balance between circuit expressivity and trainability, selecting circuits that potentially avoid both under-parameterization and the barren plateau problem associated with over-parameterization, as suggested by studies Holmes et al. (2022); Cerezo et al. (2021) on quantum circuit complexity and trainability. This screening mechanism effectively eliminates candidate circuits that clearly do not meet task requirements. To complement the limitations of pure topological analysis, QUEST-A introduces an expressivity assessment $D(c)$ based on KL divergence. This method quantifies circuit expressivity by comparing the quantum state distribution generated by the circuit with the ideal Haar random distribution Jiang (2005). Specifically, for a quantum circuit c, its expressivity can be quantified by the following
$E(c)=-D_{KL}(P(c,F)||P_{Haar}(F))$ (2)
where $P(c,F)$ is the fidelity distribution of quantum states generated by circuit c, $P_{Haar}(F)$ is the Haar random distribution, and $D_{KL}$ denotes KL divergence. $F=|\langle\psi_{\theta}|\psi_{\phi}\rangle|^{2}$ defines the fidelity of quantum states, where $\theta$ and $\phi$ are the parameters of circuit c. While the exact value of quantum state fidelity can be calculated from the density matrices before and after quantum state evolution, in practice, $P(c,F)$ is approximated by random parameter sampling of circuit $c$. The analytical form of $P_{Haar}(F)$ is:
$P_{Haar}(F)=(N-1)(1-F)^{N-2}$ (3)
where N is the dimension of the Hilbert space. A larger $E(c)$ value indicates higher expressivity, reflecting the circuit’s ability to explore a larger portion of the Hilbert space.
The KL divergence metric captures subtle circuit expressivity characteristics that might be overlooked by simple topological analysis, providing potential performance estimates, especially when considering quantum properties such as entanglement and superposition. This approach aligns with work research Sim et al. (2019) on quantum circuit expressivity, emphasizing the importance of state distribution uniformity in evaluating circuit performance. In practical work, the relationship between KL divergence and candidate circuit performance does not exhibit a simple linear correlation. Instead, it provides a range of circuits with relatively stable performance, which can be used to filter out a series of circuit structures with lower training potential.
Figure 6: Relationship between KL divergence and quantum circuit performance in task scenarios.The test evaluates KL divergence for $200$ random circuits using data re-uploading circuits as templates. These circuits were trained on an MNIST $4$-class classification task until convergence. Outliers have been removed.
By combining $P(c)$ and $E(c)$, QUEST-A achieves rapid and relatively accurate screening of high-potential candidate circuits. The integration of these proxy indicators, based on topological complexity and expressive potential reflecting the multifaceted nature of quantum circuit performance, providing effective individuals for subsequent fine-tuning optimization.
3.3 Focused Architecture Optimization
After completing the rapid architecture pruning, the QUEST-A framework enters a more refined optimization phase. The primary objective of this stage is to efficiently optimize the parameters of the selected candidate quantum circuits. Through computationally intensive methods, this phase aims to fully exploit the potential of each circuit structure.
In the field of QAS, gradient descent has been widely recognized as the most prevalent fine-grained optimization method. Given a set of candidate quantum circuits $\mathcal{P}^{\prime}_{g}=\{c_{1},c_{2},...,c_{K}\}$, where each circuit $c_{i}$ is composed of a structure $A_{i}$ and parameters $\theta_{i}$, the objective is to find the optimal parameters for each circuit as
$\theta_{i}^{*}=\arg\min_{\theta}\mathcal{L}(c_{i},\theta),$ (4)
where $\mathcal{L}(c_{i},\theta)$ is the loss function for circuit $c_{i}$. However, as the number of candidate quantum circuits increases, the computational resources required for gradient descent optimization grow exponentially.
To address this issue, parameter reuse techniques have been extensively adopted in QAS. These techniques enhance optimization efficiency by transferring and sharing training parameters among different candidate circuits. The QUEST-A combines parameter reuse with evolutionary algorithms, developing a more adaptive optimization strategy based on this foundation.
Within the evolutionary algorithm framework of QUEST-A, parameter reuse serves not merely as a means to accelerate training but as an effective mechanism for propagating the advantages of each generation of QAS. Specifically, in each generation of candidate circuits, the best-performing individual is designated as the training agent. After undergoing intensive optimization, its parameters are selectively reused by other individuals in the next generation, thus disseminating superior features throughout the entire population.
Let $c_{best}$ be the best-performing individual in the current generation as
$c_{best}=\arg\max_{c\in\mathcal{P}^{\prime}_{g}}E(c),$ (5)
where $E(c)$ is the fitness function incorporating both the DAG path count $P(c)$ and KL divergence $D(c)$:
$E(c)=f(P(c),D(c)).$ (6)
The parameter reuse strategy can be formalized as:
$\theta_{i}^{new}=\alpha\theta_{best}+(1-\alpha)\theta_{i}.$ (7)
where $\alpha\in[0,1]$ is the reuse rate, and $\theta_{best}$ are the parameters of $c_{best}$. In actual tasks, $\alpha$ is set to be very close to $1$ to balance the utilization of posteriori knowledge and the diversity of search.
The rationale behind this method is twofold: firstly, based on the characteristics of evolutionary algorithms, individuals within the same generation typically possess similar circuit structures, providing a natural foundation for parameter reuse. Secondly, the re-uploading circuit template adopted by QUEST-A enables almost lossless parameter transfer between structurally similar circuits. This method achieves efficient transmission of advantageous features at the parameter level, not only improving search efficiency but also enhancing the algorithm’s ability to explore global optima in complex search spaces.
3.4 Evolutionary Synthesis of Search Strategies
A key process of the QUEST-A framework lies in its integration of the rapid architecture pruning described in Section  3.2. and the focused architecture optimization introduced in Section  3.3 into a unified evolutionary algorithm framework. This integration adopts the principles of evolutionary algorithms to guide the QAS process, combining the discrete optimization problem of circuit structures with the continuous optimization of circuit parameters.
In this two-stage joint optimization process, rapid architecture pruning and focused architecture optimization work collaboratively, forming an efficient search mechanism. Rapid architecture pruning acts as a preliminary selection mechanism, identifying promising candidate architectures. Meanwhile, focused architecture optimization conducts a refined adaptation process, fully exploiting the potential of each candidate architecture through parameter optimization. The alternation of these two processes creates a natural and efficient evolutionary cycle, driving the entire search process to continuously evolve towards superior solutions.
In the evolutionary framework of QUEST-A, the quantum circuit structure and parameters collectively constitute the genes of individuals, providing a rich space for evolutionary operations. Rapid architecture pruning and focused architecture optimization form the environment for individual survival, with DAG path count, KL divergence, and loss function values serving as multi-level fitness evaluation indicators.
This framework achieves a combination of discrete structure search and continuous parameter optimization. Rapid architecture pruning resembles large-scale structural mutation, while focused architecture optimization corresponds to small-scale parameter fine-tuning, ensuring both the breadth of the search space and the precision of optimization. The population characteristic of the evolutionary algorithm allows QUEST-A to simultaneously maintain and evaluate multiple candidate solutions. The supercircuit structure in QUEST-A can be viewed as a dynamically evolving gene pool, realizing effective knowledge transfer between generations through the parameter reuse mechanism, accelerating the optimization process and enhancing the algorithm’s adaptability.
The adaptability of the evolutionary algorithm enables QUEST-A to dynamically adjust strategies during the search process, gradually focusing on more promising search areas while maintaining the ability to explore global optimal solutions, effectively mitigating the common problem of premature convergence in QAS. The framework of QUEST-A is described as follows:
Algorithm 1 QUEST-A Framework
1:  Initialize supercircuit $S$ with architecture $A$ and parameters $\theta$
2:  Define $E(c)=f(P(c),D(c))$Fitness function}
3:  for generation $g=1$ to $G$ do
4:     $\mathcal{P}_{g}\leftarrow$ Sample from $S$
5:     for $c\in\mathcal{P}_{g}$ do
6:        Evaluate $P(c)$DAG path count}
7:        Evaluate $D(c)$KL divergence}
8:     end for
9:     $\mathcal{P}^{\prime}_{g}\leftarrow$ Select top $K$ candidates based on $E(c)$
10:     $c_{best}\leftarrow\arg\max_{c\in\mathcal{P}^{\prime}_{g}}E(c)$
11:     for $c_{i}\in\mathcal{P}^{\prime}_{g}$ do
12:        $\theta_{i}^{*}\leftarrow\arg\min_{\theta}\mathcal{L}(c_{i},\theta)$Gradient descent}
13:        $\theta_{i}^{new}\leftarrow\alpha\theta_{best}+(1-\alpha)\theta_{i}^{*}$Parameter reuse}
14:     end for
15:     Update $S$ with knowledge from $c_{best}$
16:     $\mathcal{P}_{g+1}\leftarrow$ Evolve($\mathcal{P}^{\prime}_{g}$) {Crossover and mutation}
17:     Update $A$ and $\theta$ in $S$ based on $\mathcal{P}_{g+1}$
18:  end for
19:  return  $\arg\max_{c\in\mathcal{P}_{G}}E(c)$
By leveraging both coarse-grained pruning and fine-grained optimization within an evolutionary framework, QUEST-A aims to overcome the challenges of computational complexity and premature convergence often encountered in QAS. To validate the effectiveness and efficiency of this approach, we now turn to a comprehensive experimental analysis across various quantum computing tasks.
4 Experiment Results and Performance Analysis
4.1 Experiment Framework
A series of experiments were designed to validate the effectiveness and performance advantages of the QUEST-A method. These experiments include: signal representation tasks to demonstrate QUEST-A’s performance in generative tasks, particularly in handling complex signals; image classification tasks to verify QUEST-A’s performance in practical application scenarios; and variational quantum eigensolver (VQE) tasks to showcase its potential in quantum advantage domains. In each experiment, models were trained five times, with the best performance selected as the comparison result. Detailed information for experiments can be found in  B.
Numerical simulation experiments were conducted on a Python-based simulation platform, utilizing PennyLane Bergholm et al. (2018) , Mindspore Quantum Xu et al. (2024), and TensorFlow Quantum Broughton et al. (2020). Experiments in a real quantum computing environment were performed on the ibm_kyiv backend.
4.2 Task I: QUEST-A in Insignal Representation
The insignal representation task aims to learn the intrinsic representation of signals through modeling, with its core objective being the learning of a function that maps input coordinates to corresponding signal values. This task can be viewed as a specialized form of continuous function approximation or signal reconstruction. In the field of quantum computing, this task demonstrates the capability of quantum models in continuous function approximation and also showcases potential in implicit neural representations tasks Mildenhall et al. (2021).
This experiment utilizes the same audio and image datasets as the benchmark models Zhao et al. (2024), ensuring consistency in data sources and preprocessing methods. Specifically, for the image representation task, images of astronaut, camera, and coffee from the scikit-image dataset Van der Walt et al. (2014) were selected.These images were first downsampled to a resolution of $32×32$ pixels and then resized through a linear layer before being input into a 10-qubit quantum model. The model processed the data through three iterations of data re-uploading for learning. Additionally, the model is equipped with two variational layers, structurally identical to the re-uploading circuits, both at the initial stage and prior to the measurement stage. These layers serve to optimize the initial state of quantum evolution and the basis vectors for measurement.
The evolutionary algorithm was configured to run for 10 generations, with each generation comprising 10 individuals. For the dataset $D={(x_{i},y_{i})}$, where $x_{i}$ represents the coordinates of the signal $y_{i}$, the mean squared error (MSE) is utilized as the performance metric for the model output $y_{i}\prime$. The fitness function during the evolutionary process is defined as follow
$Fitness:=-\log\left(\frac{1}{|D|}\sum_{i=1}^{|D|}(y_{i}-y_{i}^{\prime})^{2}\right).$ (8)
The learning progression and final results of the model on the Astronaut, $\textit{Camera},and$Coffee images are visually represented in Fig. 7.
Figure 7: Performance of QUEST-A in image representation tasks. The numerical values beneath each image represent the corresponding MSE. The red area illustrates the evolution of MSE for the best-performing individual in QUEST-A over generations. The green area shows the original images as the task benchmark. The blue area presents the results and MSE values for baseline Zhao et al. (2024); Sitzmann et al. (2020).
The results demonstrate that in the initial stages of the QUEST-A method’s search process, the candidate circuit population obtained through rapid screening performed poorly, capturing almost no effective information. However, as iterations progressed, the performance of candidate circuits showed a rapid improvement trend. For the best-performing image, QUEST-A reduced the MSE from $4.0\times 10^{-3}$ of the baseline model to $2.2\times 10^{-3}$, achieving approximately a $45\%$ reduction in error rate. In the Astronaut image, which exhibits a more complex frequency pattern, QUEST-A similarly achieved a low error rate, demonstrating visually discernible advantages in capturing image details. Compared to the two baseline methods, QUEST-A exhibited superior performance in processing high-frequency information relative to the SOTA quantum methods, resulting in more precise retention of details and edge information in the generated images. The performance of QUEST-A stems from its further enhancement of the inherent advantages of data re-uploading quantum circuits in continuous function approximation.
To further validate QUEST-A’s capability in processing different types of signals, we also conducted experiments on audio representation tasks, with the results presented in Fig. 8.
Figure 8: Performance of QUEST-A in audio representation tasks. The red area illustrates the performance evolution of QUEST-A during the search process and its final results. The green area shows the original audio signals as the task benchmark. The blue area presents the performance of two baseline models on this task.
The results of the audio representation task further confirm the advantages of the QUEST-A model in signal recognition. Although in terms of MSE, the QUEST-A model achieved only a $7.3\%$ reduction in error rate compared to the baseline model, its performance in audio amplitude details processing surpassed state-of-the-art quantum methods. From the perspective of frequency graphs, the results generated by QUEST-A more accurately reproduced the frequency patterns of the benchmark audio, especially demonstrating more refined reproduction in the high-frequency components. While these details may contribute minimally to the overall MSE calculation, they are crucial for the perception of audio quality.
During the training process, we observed that the QUEST-A model quickly learned low-frequency patterns in the initial stages, similar to its performance in the image representation task. In the subsequent training process, the model made fine adjustments to the high-frequency components of the waveform and rapidly converged. This learning pattern further demonstrates QUEST-A’s adaptability in handling different types of insignal.
Table 1: Performance of QUEST-A and SOTA models in signal representation tasks.
Method Sound Representation Image Representation
Cello params Astronaut Camera Coffee paramsa
SIREN 5.5b 691 9.0 1.5 2.3 701
QIREN 5.5 649 4.0 1.1 1.5 657
[…]
  • a
    The params column indicates the number of parameters used by each model.
  • b
    Performance is measured by the MSE(×$\times 10^{-3}$) between the representation results and the task benchmark, lower MSE values indicate better performance.
4.3 Task II: QUEST-A in Image Classification
Unlike signal representation tasks that focus on function fitting, image classification tasks emphasize the model’s understanding of data distribution and its ability to construct decision boundaries. This study uses the widely adopted MNIST and Fashion-MNIST datasets as benchmarks, comparing QUEST-A with several state-of-the-art quantum baseline methods, including QUILT Silver et al. (2022), Élivágar Anagolum et al. (2024), and QNAS Wang et al. (2022). To thoroughly evaluate QUEST-A’s performance in the NISQ era, we designed three experimental scenarios: numerical simulation, noise simulation, and experiments under real noisy environments. For the simulated noisy environment, we employed a composite noise model incorporating bit flip, phase flip, and depolarizing channel noise.
In terms of preprocessing, all $28\times 28$ pixel images first undergo principal component analysis (PCA), after which the processed data is input into a 10-qubit quantum circuit. The model performs four iterations of data re-uploading and uses cross-entropy as the loss function. The experimental results and performance comparisons on the MNIST and Fashion-MNIST datasets are presented in Tab. 2.
Table 2: Performance of QUEST-A and SOTA models in classification tasks.
Task NNSa SNS RNE
Quilt QNAS QUEST-A (ours) Élivágarb QUEST-A (ours) QNAS Élivágar QUEST-A (ours)
MNIST-2 94.5c 95.0 99.9 97.0 99.7 95.0 93.0 98.5
MNIST-4 80.8 77.0 97.7 64.0 90.5 71.0 65.0 92.0
FMNIST-2 - - 98.1 97.0 95.8 90.0 97.0 94.3
FMNIST-4 77.7 - 91.2 76.0 86.0 85.0 73.0 82.4
  • a
    The table presents performance under three conditions: noiseless numerical simulation (NNS), simulated noisy simulation (SNS), and real noisy environment (RNE).
  • b
    For baseline works, if multiple results were available under different conditions (e.g. different basic gate sets or noise environments), the best performance was selected as the final performance for the baseline model.
  • c
[…]
QUEST-A demonstrated significant performance and robustness on both MNIST and Fashion-MNIST datasets. Compared to baseline models, QUEST-A achieved higher classification accuracy in noiseless environments (NNS). Notably, when noise was introduced (SNS), QUEST-A’s performance showed only a slight decrease. This noise robustness became more evident as task difficulty increased; when transitioning from 2-class to 4-class classification, QUEST-A’s performance degradation was less pronounced than that of baseline models. QUEST-A performed well on both MNIST and Fashion-MNIST datasets, indicating its generalization capability and adaptability. Further validation of QUEST-A’s performance in classification tasks is presented as Tab. 3 and Fig. 9.
Table 3: Performance of QUEST-A in multi-class classification tasks.
Dataset Number of Classesa
2 4 6 8 10
MNIST 0.9994 (±$(\pm 0.3{\times}10^{-3})$ 0.9736 (±$(\pm 3.2{\times}10^{-3})$ 0.9391 (±$(\pm 3.5{\times}10^{-3})$ 0.9129 (±$(\pm 1.7{\times}10^{-3})$ 0.8545 (±$(\pm 2.0{\times}10^{-3})$
FMNIST 0.9810 (±$(\pm 1.3{\times}10^{-3})$ 0.9087 (±$(\pm 4.8{\times}10^{-3})$ 0.8513 (±$(\pm 2.7{\times}10^{-3})$ 0.7576 (±$(\pm 3.5{\times}10^{-3})$ 0.7636 (±$(\pm 2.6{\times}10^{-3})$
  • a
    Performance of QUEST-A on MNIST and Fashion-MNIST for multi-class classification tasks (±($\pm$( ± standard deviation))))
Figure 9: Heatmap of classification accuracy vs. number of re-uploading.
In-depth analysis of image multi-classification tasks further validates the robustness of the QUEST-A model. QUEST-A demonstrated high performance stability across classification tasks ranging from $2$ to $10$ classes. Although model performance showed an expected declining trend as the number of classes increased, QUEST-A still maintained a high accuracy of over $75\%$ in the most challenging 10-class classification task. A noteworthy phenomenon is that increasing the number of trainable layers did not bring about the expected significant performance improvement. On the contrary, in some cases, a slight performance decline was observed. This unexpected result may stem from two factors: firstly, the model may have experienced a certain degree of overfitting in determining decision boundaries; secondly, the current task settings may not have fully utilized the potential of the model’s assumptions.
Above experimental results not only further confirm QUEST-A’s excellent robustness in handling multi-classification tasks but also provide insights into the model’s potential performance in more complex application scenarios.
4.4 Task III: QUEST-A in Variational Quantum Eigensolver
The VQE task aims to estimate the ground state energy of a given Hamiltonian. In the quantum computing framework, this involves optimizing a parameterized quantum circuit to minimize the expected energy of the system. The VQE task primarily examines the model’s ability to prepare complex quantum states, which should closely approximate the ground state of the target Hamiltonian, while also evaluating whether the ansatz design has sufficient expressive power to represent the solution space of the problem.
This experiment selects three typical systems as VQE objects: the $H_{2}$ molecule (4 qubits), the Heisenberg model (5 qubits), and the Transverse Field Ising Model (TFIM, 6 qubits). Due to the unique characteristics of the VQE, QUEST does not employ data re-uploading circuits in this task. The gate set used for the task includes RX, RY, RZ, and CNOT gates. The Hamiltonian for the $H_{2}$ molecule follows the construction method from previous work Kandala et al. (2017). The Hamiltonians for TFIM and the Heisenberg model are defined as follows
$H_{TFIM}=\sum_{i=1}^{n}Z_{i}Z_{i+1}-X_{i},$ (9)
$H_{Heis}=\sum_{i=1}^{n}X_{i}X_{i+1}+Y_{i}Y_{i+1}+Z_{i}Z_{i+1}+\sum_{i=1}^{n}z_{i},$ (10)
where $n$ denotes the number of qubits. The model’s performance is measured by the difference between the estimated value $E$ and the theoretically exact ground state energy $E_{0}$, i.e., $E-E_{0}$. To ensure fair comparison, QUEST-A and the baseline models use the same number of qubits and similar parameter counts. In each experiment, the models are run multiple times, with the best performance selected as the final result.
We selected two representative baseline models for comparison, TF-QAS He et al. (2024b) and PQAS Zhang et al. (2021), as fast architecture search method. Tab. 4. presents the performance of QUEST-A and these baseline models on three quantum systems along with detailed comparison results.
Table 4: Performance of QUEST-A and SOTA models in VQE tasks.
Systems Qubit PQAS TF-QAS QUEST-A (ours)
[…]
Heisenberg 5 $8.9\times 10^{-2}$ $1.2\times 10^{-3}$ $1.7\times 10^{-14}$
TFIM 6 $9.0\times 10^{-2}$ $1.2\times 10^{-5}$ $2.8\times 10^{-8}$
  • a
    Model performance is defined by the difference between the model’s estimated value and the theoretical exact value.
QUEST-A achieved the highest precision in all test cases, demonstrating an advantage of up to 11 orders of magnitude over state-of-the-art methods. Compared to similar fast architecture search methods, QUEST-A not only successfully completed the ground state energy prediction task within chemical accuracy $(1.6\times 10^{-3}Ha)$ but also achieved more precise predictions. As the system size increased from $4$ qubits to $6$ qubits, QUEST-A maintained its performance advantage. This outcome reflects QUEST-A’s ability in preparing complex quantum states and approximating the ground state of target Hamiltonians. Its exceptional performance may stem from QUEST-A’s ability to avoid getting trapped in local optima during the search process, as well as its enhanced sensitivity to the fine structure of quantum states. Furthermore, the fact that this task did not utilize existing high-performance ansatzes corroborates that QUEST-A does not rely on specific circuit patterns, but rather can adaptively discover quantum circuit structures suitable for specific problems. This enhanced expressive capability gives QUEST-A potential advantages in handling more complex quantum systems.
In conclusion, we evaluated QUEST-A across three distinct tasks: insignal representation, image classification, and VQE. In the insignal representation task, QUEST-A achieved up to $81.8\%$ performance improvement compared to SOTA methods, with a slightly lower parameter count. For image classification, it demonstrated superior accuracy and robustness, maintaining over $75\%$ accuracy in $10$-class tasks and showing adaptability to noisy environments. In VQE tasks, QUEST-A outperformed existing methods by up to $11$ orders of magnitude in precision. These results consistently demonstrate QUEST-A’s performance and adaptability in handling complex quantum tasks, validating its effectiveness as a QAS method.
5 Discussion
The QUEST-A framework demonstrates how quantum architecture search can effectively bridge quantum and classical computing paradigms. While quantum computation excels at rapid estimation and parallel processing (as shown in KL divergence calculation), classical optimization techniques provide reliable solutions for parameter tuning and precise optimization. This complementary approach provides guiding principles for developing hybrid algorithms. Another exploration reveals that QAS can be utilized to optimize not only intermediate quantum circuits but also data encoding and measurement schemes. Recent works Wang et al. (2023); Thanos et al. (2024) in automated circuit design, combined with advances in quantum embedding Havlíček et al. (2019) and measurement Quek et al. (2021), indicate promising directions for developing task-specific quantum architectures.
Furthermore, the flexibility of the QUEST-A framework in the rapid screening phase allows for easy incorporation of hardware-specific factors. However, caution must be exercised when considering hardware noise in QAS. Although many current works Wang et al. (2022); Anagolum et al. (2024); Lu et al. (2023); He et al. (2024a) consider quantum hardware characteristics such as qubit architecture and noise properties, the time-varying nature of quantum noise may quickly render static noise models obsolete, leading to a loss of circuit adaptability. Therefore, future research needs to develop more dynamic and adaptive circuit noise mitigation strategies to better address noise challenges in practical quantum systems.
6 Conclusion
In this work, we propose QUEST-A, a QAS framework that unifies discrete pruning and continuous optimization processes. This unification is achieved by decomposing the QAS problem into two subproblems: optimal circuit structure retrieval and parameter optimization. The framework comprises three core components: rapid architecture pruning mechanism, focused architecture optimization strategy, and evolutionary algorithm framework. Extensive experiments demonstrate QUEST-A’s superior performance across signal representation, image classification, and VQE tasks, validating its enhanced expressivity and precision.
Our exploration reveals that QAS requires a synergy of discrete search and continuous optimization paradigms, while leveraging validated ansatze as prior knowledge. The essence of QAS lies in its knowledge-driven dynamic nature, where the process accumulates understanding of advantageous features during circuit discovery, analogous to knowledge distillation and transfer learning in modern deep learning.
Future work will build upon these understandings of the QAS problem to design more efficient and precise search methods from the perspective of quantum circuit expressibility.
Authors’ contributions
Lian-Hui Yu: Methodology, Software. Xiao-Yu Li: Conceptualization, Project administration. Geng Chen: Writing - Original Draft, Visualization. Qin-Sheng Zhu: Validation. Hui Li: Formal analysis. Guo-Wu Yang: Supervision.
Availability of supporting data
All data generated or analysed during this study are available and included in this published article. Code for our work has been open-sourced.
Acknowledgements
This work was supported by the Open Fund of Advanced Cryptography and System Security Key Laboratory of Sichuan Province (Grant No. SKLACSS-202210), National Natural Science Foundation of China (Grant No. 62472072 and No. 62172075).
Appendix A Preliminaries of Quantum Circuit
The essence of quantum computing lies in embedding data into Hilbert space and processing information using quantum properties such as superposition and entanglement. Within this framework, the role of quantum basic gates is to ensure that the operable domain fully covers the target Hilbert space, enabling precise control and manipulation of quantum states.
In QAS task, basic gates are the fundamental units of circuit design. The main task of QAS is to determine the position and parameters of these basic gates to construct the optimal quantum circuit. The basic gates used in this paper include $R_{X}$, $R_{Y}$, $R_{Z}$, and $CNOT$ gates, whose mathematical forms are as follows
$RX(\theta)=\begin{pmatrix}\cos(\theta/2)&-i\sin(\theta/2)\\ -i\sin(\theta/2)&\cos(\theta/2)\end{pmatrix},RY(\theta)=\begin{pmatrix}\cos(\theta/2)&-\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{pmatrix}$ (11)
$RZ(\theta)=\begin{pmatrix}e^{-i\theta/2}&0\\ 0&e^{i\theta/2}\end{pmatrix},CNOT=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (12)
where, CNOT is a two-qubit gate, while the others act on single qubits. As introduced in Section  2.1, once the quantum circuit structure is determined, the variational quantum circuit can be trained through a fixed process.
It is worth noting that whether to use classical gradient optimization methods to guide PQC learning remains an open question. For instance, the promising physical neural networks (PNN) Momeni et al. (2023) approach adaptively circumvents gradient-based optimization processes.
Appendix B Experimental details
The experimental settings for our work are presented in the Tab. 5
Table 5: Detailed parameter settings for all experiments in our work
Parameter Signal Representation Image Classification VQE
Number of qubits 8 10 5 ($Heisenberg$) 4 ($H_{2}$), 6 ($TFIM$)
Data re-uploading 2 3 -a
Number of generations 10 10 10
Population size 5 5 5
Loss metric MSE Crossentropy -
Trainable parametersb 601 ($Sound$) 609 ($Image$) 175 175 ($Heisenberg$) 105 ($H_{2}$), 420 ($TFIM$)
  • a
    The VQE task does not utilize specific circuit templates, nor does it employ gradient-based optimization processes.
  • b
    Experimental parameters are set to be comparable with baseline models.
In the search process, the best-performing individual is selected as the final performance metric for the experiment. All experiments were conducted independently five times, with the best performance across these runs reported as the final result. The code for QUEST-A is available.
Appendix C Ablation Analysis
We conducted a set of ablation experiments using the Fashion-MNIST 10-class classification task as the experimental scenario. The experiments investigated the following three cases: (i) Training random circuits. (ii) Training after DAG screening only; (iii) Training after both DAG and KL screening;
For random circuit training (case i), we used $5$ randomly generated circuits for training. For the rapid screening cases (cases ii and iii), we screened from $100$ candidate circuits and selected the top $5$ optimal circuits. All experiments were terminated upon convergence of the loss function.
Figure 10: Loss trend over epochs for three case. Training progress shown until convergence of case i.
Circuits screened by both DAG and KL methods (case iii) exhibited faster convergence and superior performance compared to the other two cases. Random circuits (case i) demonstrated slower convergence and inferior final performance relative to the screened circuits. Comparing random circuits with those that underwent rapid screening, it is evident that the screening process effectively altered the model’s initial performance and convergence capability, indicating the selection of circuits with higher performance potential. Moreover, the more refined screening process (case iii versus case ii) yielded further improvements in both convergence speed and final performance.